# HippoMM: Hippocampal-inspired Multimodal Memory for Long Audiovisual Event Understanding

### v1

@@ A.6.4 Semantic/Temporal Reasoning (TS)
Example: Semantic/Temporal Reasoning Question: “What was one significant challenge the main character faced during the experiment?” Options: A: Finding time to eat breakfast B: Dealing with brain fog C: Avoiding social media D: Lack of exercise options Explanation: The main character starts the day with a homemade green juice made from various fruits and vegetables.
Appendix B HippoMM Implementation Details
This appendix provides the specific implementation choices, models, and parameters used in the HippoMM system, complementing the conceptual description in Section 4.
@@ Appendix B HippoMM Implementation Details
B.1 Memory Formation Implementation
Temporal Pattern Separation (Sec. 4.2.1)
@@ B.1 Memory Formation Implementation
  • Segment Duration: The resulting segments are constrained to be between $t_{min}=5$ seconds and $t_{max}=10$ seconds long. Shorter segments following boundary detection are merged with preceding ones, and longer intervals between boundaries are split to adhere to this constraint, defining the temporal windows $[t_{start,i},t_{end,i}]$.
Perceptual Encoding (Sec. 4.2.2)
Each segment $i$ (defined by $[t_{start,i},t_{end,i}]$) is processed to extract multimodal features:
@@
  • ShortTermMemory Object ($m_{i}$): These processed elements are aggregated and stored in a ShortTermMemory object $m_{i}$, as defined in Eq. 7. It encapsulates: pointers to the raw audiovisual snippets ($\mathbf{C}_{i}$), the sequence of ImageBind embeddings ($\mathbf{E}_{i}$), the aggregated Whisper transcriptions and Qwen-VL visual descriptions ($\mathbf{T}_{i}$), segment start/end timestamps ($t_{i},s_{i}$), and related metadata.
Memory Consolidation (Sec. 4.2.3)
@@
B.2 Memory Retrieval Implementation
Query Analysis (Sec. 4.3.1)
@@ B.2 Memory Retrieval Implementation
  • Fast Retrieval ($\Phi_{\text{fast}}$): Retrieval is first attempted using the abstract textual summaries ($\mathbf{S}_{\theta_{k}}$) stored in the collection of ThetaEvent objects (=$\mathcal{E}=\{\theta_{k}\}$). Confidence scores are determined by internal LLM calibration (Qwen2.5-VL, see Table 5) on a scale of 0-1.
  • Detailed Recall ($\Psi_{\text{detailed}}$): If the confidence score from Fast Retrieval falls below a threshold $\tau=0.75$, retrieval escalates to the Detailed Recall pathway. This pathway searches the consolidated ShortTermMemory objects (=$\mathcal{M}=\{m_{k}\mid k\in K\}$), guided by the query type determined by $\mathcal{Q}_{\text{type}}(q)$.
Detailed Recall Pathways (Sec. 4.3.2)
@@
  • Similarity Search (Feature/Semantic): For single-modality or cue-based retrieval (Visual or Auditory query type), cosine similarity is used to find the top-$k$ ShortTermMemory segments ($m_{k}$) whose representative embeddings ($\mathbf{v}_{k}$) are most similar to $q_{embed}$ (Feature Search), with $k=5$. Text-based search over the aggregated descriptions/transcriptions ($\mathbf{T}_{k}$) serves as a fallback or alternative (Semantic Search). See Tables 6 and 7 for related selection prompts.
@@
Appendix C Additional Experimental Details
C.1 Ablation Studies Parameter Settings
Acronyms refer to components described in Sec. 4: FR (Fast Retrieval, $\Phi_{\text{fast}}$), DR (Detailed Recall, $\Psi_{\text{detailed}}$), AR (Adaptive Reasoning, $\Gamma$).
@@ C.1 Ablation Studies Parameter Settings
C.2 Performance Metrics
  • Processing Time (PT): Total time to execute the Memory Formation phase (Sec. 4.2) on raw videos.
  • Average Response Time (ART): Average time to execute the Memory Retrieval phase (Sec. 4.3) and generate an answer $a$ for a query $q$.
@@ C.3 Generalization Evaluation
C.4 Additional Implementation Parameters
@@
Task Description Please provide a concise one sentence summary of this event based on the video frames descriptions and audio transcription. What is happening in this event? Input Format Image descriptions: [Concatenated frame captions from key visual context $\mathbf{I}_{\text{visual}}(m_{k})$] Audio transcription: [Concatenated transcribed text from key auditory context $\mathbf{A}_{\text{audio}}(m_{k})$] Output Requirements One sentence summary ($\mathbf{S}_{\theta_{k}}$) capturing the integrated multimodal essence of the event Focus on key actions, subjects, and setting Maintain factual accuracy based on provided descriptions Avoid introducing information not supported by the input
@@
Task Description Classify this question into one of these categories based on what type of information is needed to answer it: Classification Categories VIDEO - Questions specifically about visual elements, appearances, or actions that need frame-by-frame analysis, e.g., “what is the main character holding?” (Corresponds to Visual $V$) AUDIO - Questions about sounds, speech, or audio content that need audio analysis, e.g., “what does the main character say/mention?” (Corresponds to Auditory $A$) VIDEO+AUDIO - Questions that require associating both visual and audio information, e.g., “what is the main character doing while saying/mentioning something?” (Corresponds to Cross-modal $V\times A$) SUMMARY - Questions that focus on the overall content or gist of the video, e.g., “what is the main character doing?” (Corresponds to Semantic/Gist) Guidelines If question is about visual details, appearances, or actions, classify as VIDEO If question is about sounds, speech, or audio content, classify as AUDIO If question requires both visual and audio information, classify as VIDEO+AUDIO If question is about the overall content of the video, classify as SUMMARY Input Question: [user query $q$] Output Format Return ONLY one of these exact words: VIDEO, AUDIO, VIDEO+AUDIO, SUMMARY
@@
Task Description Given the following question and relevant event summaries ($\mathbf{S}_{\theta_{k}}$ from $\mathcal{E}$), analyze whether the question can be answered directly or needs specific analysis (triggering Detailed Recall $\Psi_{\text{detailed}}$). Input Question: [user query $q$] Event Summaries: [List of relevant event summaries $\mathbf{S}_{\theta_{k}}$ retrieved based on $q$] Guidelines for analysis General questions about overall video content (likely classified as SUMMARY) should be attempted directly from summaries. Questions about specific visual details (VIDEO), sounds/speech (AUDIO), or combined elements (VIDEO+AUDIO) often require specific analysis beyond summaries. Estimate confidence based on whether summaries contain sufficient information. Output Format Must be in one of these two structures: 1. If answerable from summaries:
ANSWER: <your detailed answer based ONLY on summaries>
@@ 2. If requiring specific analysis (summaries insufficient):
Table 6: Visual Segment Selection Prompt (Part of $\Psi_{\text{detailed}}$ Semantic Search)
Task Description Given a question (potentially classified as VIDEO or VIDEO+AUDIO), descriptions from consolidated memory segments ($\mathbf{T}_{k}$), and an optional element to search for, identify at most 5 relevant frame indices (within segments $m_{k}$) for answering the question. Input Question: [user query $q$] Element to search for: [search query derived from $q$] Frame descriptions (from relevant $\mathbf{T}_{k}$): [Numbered list of frame descriptions/captions] Instructions Return ONLY numbers separated by commas (e.g., “0,3,5,8,12”) corresponding to indices within the provided list. Return at most 5 indices. Do not include any other text, explanations, or spaces. If fewer than 5 frames are relevant, return fewer indices. Example good responses “0,3,5,8,12” “1,4,7” Output Format Your response (numbers only, comma-separated):
Table 7: Audio Segment Selection Prompt (Part of $\Psi_{\text{detailed}}$ Semantic Search)
Task Description Given a question (potentially classified as AUDIO or VIDEO+AUDIO) and transcriptions with timestamps from consolidated memory segments ($\mathbf{T}_{k}$), identify the most relevant time frames where the answer might be found. Input Question: [user query $q$] Transcriptions (with timestamps from relevant $\mathbf{T}_{k}$): [List of timestamped transcriptions] Instructions Analyze the transcriptions and identify segments most likely to contain the answer. Return a JSON array of time frames (relative to the source video). Each frame defines a start and end time.
@@
Return at most 5 time frames. Include a small buffer around identified time frames (e.g., $\pm$2 seconds, related to $\delta$). If no relevant segments found, return “[]”.
@@
Task Description You potentially have two candidate answers to the same question $q$. One derived from Fast Retrieval ($\Phi_{\text{fast}}$, using summaries $\mathcal{E}$) and another from Detailed Recall ($\Psi_{\text{detailed}}$, using detailed segments $\mathcal{M}$). Compare them and provide a final, reconciled answer $a$. Input Question: [user query $q$] Answer from summaries (if $\Phi_{\text{fast}}$ produced one): [direct_answer] Confidence from summaries: [direct_confidence] Answer from detailed analysis (if $\Psi_{\text{detailed}}$ was run): [detailed_answer] Supporting Context ($\mathcal{E}_{\text{context}}$): Sample frame captions: [Sample relevant frame captions from $r_{\text{retrieved}}$] Sample transcriptions: [Sample relevant transcriptions from $r_{\text{retrieved}}$] Instructions Compare both answers for consistency, detail, and plausibility given the context. If they agree, prefer the more detailed/confident answer. If they disagree, evaluate which answer is better supported by the specific context (captions, transcriptions). Prioritize the answer derived from detailed analysis ($\Psi_{\text{detailed}}$) if confidence in the summary-based answer ($\Phi_{\text{fast}}$) was low or if the detailed analysis provides clearly contradictory evidence. Synthesize the best possible final answer based on this comparison. For multiple choice questions, select only one final answer option. Estimate the confidence in the final reconciled answer. Output Format
@@
Task Description Based on the following retrieved context ($r_{\text{retrieved}}$) and extracted evidence ($\mathcal{E}_{\text{context}}$), please synthesize a final answer ($a$) to the question ($q$). Input Question: [user query $q$] Overall Video Context (e.g., relevant summaries $\mathbf{S}_{\theta_{k}}$ if available): [List of relevant high-level context] Relevant Retrieved Content (details from $\Phi_{\text{fast}}$ or $\Psi_{\text{detailed}}$): [List of retrieved content descriptions, e.g., frame captions, transcriptions, details from $m_{k}$] Instructions Please provide a clear and specific answer based on the integration of the overall context and the specific retrieved content. If the information is insufficient or contradictory, state that or provide the best possible reasoned answer based on the evidence. Output should be one letter if the original question was multiple choice. Output Format Answer: <Synthesized answer $a$>

### v2

@@ B.6.4 Semantic/Temporal Reasoning (TS)
Example: Semantic/Temporal Reasoning
Question: “What was one significant challenge the main character faced during the experiment?”
Options:
  • A: Finding time to eat breakfast
  • B: Dealing with brain fog
  • C: Avoiding social media
  • D: Lack of exercise options
Explanation: Later in the vlog, after the experiment has been underway for several hours, the main character states they are “dealing with a lot of brain fog,” which is affecting their focus on the task.
Appendix C HippoMM Implementation Details
This appendix provides the specific implementation choices, models, and parameters used in the HippoMM system, complementing the conceptual description in the Methodology section.
@@ Appendix C HippoMM Implementation Details
C.1 Memory Formation Implementation
Temporal Pattern Separation
@@ C.1 Memory Formation Implementation
  • Segment Duration: The resulting segments are constrained to be between $t_{min}=5$ seconds and $t_{max}=10$ seconds long. Shorter segments following boundary detection are merged with preceding ones, and longer intervals between boundaries are split to adhere to this constraint, defining the temporal windows $t_{start,i},t_{end,i}]$.
Perceptual Encoding
Each segment $i$ (defined by $t_{start,i},t_{end,i}]$) is processed to extract multimodal features:
@@
  • ShortTermMemory Object ($m_{i}$): These processed elements are aggregated and stored in a ShortTermMemory object $m_{i}$, as defined in Eq. (4). It encapsulates: pointers to the raw audiovisual snippets ($\mathbf{C}_{i}$), the sequence of ImageBind embeddings ($\mathbf{E}_{i}$), the aggregated Whisper transcriptions and Qwen-VL visual descriptions ($\mathbf{T}_{i}$), segment start/end timestamps ($t_{s,i},t_{e,i}$), and related metadata.
Memory Consolidation
@@
C.2 Memory Retrieval Implementation
Query Analysis
@@ C.2 Memory Retrieval Implementation
  • Fast Retrieval ($\Phi_{\text{fast}}$): Retrieval is first attempted using the abstract textual summaries ($\mathbf{S}_{\theta_{k}}$) stored in the collection of ThetaEvent objects ($\mathbf{\Theta}=\{\theta_{k}\}$). Confidence scores are determined by internal LLM calibration (Qwen2.5-VL, see Table 7) on a scale of 0-1.
  • Detailed Recall ($\Psi_{\text{detailed}}$): If the confidence score from Fast Retrieval falls below a threshold $\tau=0.75$, retrieval escalates to the Detailed Recall pathway. This pathway searches the consolidated ShortTermMemory objects ($\mathcal{M}=\{m_{k}\mid k\in K\}$), guided by the query type determined by $\mathcal{Q}_{\text{type}}(q)$.
Detailed Recall Pathways
@@
  • Similarity Search (Feature/Semantic): For single-modality or cue-based retrieval (Visual or Auditory query type), cosine similarity is used to find the top-$k$ ShortTermMemory segments ($m_{k}$) whose representative embeddings ($\mathbf{v}_{k}$) are most similar to $q_{embed}$ (Feature Search), with $k=5$. Text-based search over the aggregated descriptions/transcriptions ($\mathbf{T}_{k}$) serves as a fallback or alternative (Semantic Search). See Tables 8 and 9 for related selection prompts.
@@
Appendix D Cross-Benchmark Generalization
Table 4: Cross-benchmark generalization comparison. Video-MME and LongVideoBench test general comprehension; MLVU-NQA tests precise temporal retrieval, where HippoMM’s episodic memory excels. Best results in bold.
Method Video-MME LongVideoBench MLVU-NQA
InternVL2 [4] 61.2% 59.3% 48.3%
VideoLLaMA 3 [41] 66.2% 59.8% 68.3%
GPT-4o [15] 71.9% 66.7% 64.8%
HippoMM (Ours) 61.8% 55.2% 73.1%
Appendix E Additional Experimental Details
E.1 Ablation Studies Parameter Settings
Acronyms refer to components described in the Methodology section: FR (Fast Retrieval, $\Phi_{\text{fast}}$), DR (Detailed Recall, $\Psi_{\text{detailed}}$), AR (Adaptive Reasoning, $\Gamma$).
@@ E.1 Ablation Studies Parameter Settings
E.2 Performance Metrics
  • Processing Time (PT): Total time to execute the Memory Formation phase on raw videos.
  • Average Response Time (ART): Average time to execute the Memory Retrieval phase and generate an answer $a$ for a query $q$.
@@ E.3 Generalization Evaluation
E.4 Additional Implementation Parameters
@@
Task Description Please provide a concise one sentence summary of this event based on the video frames descriptions and audio transcription. What is happening in this event?
Input Format
  • Image descriptions: [Concatenated frame captions from key visual context $\mathbf{I}_{\text{visual}}(m_{k})$]
  • Audio transcription: [Concatenated transcribed text from key auditory context $\mathbf{A}_{\text{audio}}(m_{k})$]
Output Requirements
  • One sentence summary ($\mathbf{S}_{\theta_{k}}$) capturing the integrated multimodal essence of the event
  • Focus on key actions, subjects, and setting
  • Maintain factual accuracy based on provided descriptions
  • Avoid introducing information not supported by the input
@@
Task Description Classify this question into one of these categories based on what type of information is needed to answer it:
Classification Categories
  • VIDEO - Questions specifically about visual elements, appearances, or actions that need frame-by-frame analysis, e.g., “what is the main character holding?” (Corresponds to Visual $V$)
  • AUDIO - Questions about sounds, speech, or audio content that need audio analysis, e.g., “what does the main character say/mention?” (Corresponds to Auditory $A$)
  • VIDEO+AUDIO - Questions that require associating both visual and audio information, e.g., “what is the main character doing while saying/mentioning something?” (Corresponds to Cross-modal $V\times A$)
  • SUMMARY - Questions that focus on the overall content or gist of the video, e.g., “what is the main character doing?” (Corresponds to Semantic/Gist)
Guidelines
  • If question is about visual details, appearances, or actions, classify as VIDEO
  • If question is about sounds, speech, or audio content, classify as AUDIO
  • If question requires both visual and audio information, classify as VIDEO+AUDIO
  • If question is about the overall content of the video, classify as SUMMARY
Input Question: [user query $q$]
Output Format Return ONLY one of these exact words: VIDEO, AUDIO, VIDEO+AUDIO, SUMMARY
@@
Task Description Given the following question and relevant event summaries ($\mathbf{S}_{\theta_{k}}$ from $\mathbf{\Theta}$), analyze whether the question can be answered directly or needs specific analysis (triggering Detailed Recall $\Psi_{\text{detailed}}$).
Input Question: [user query $q$]
Event Summaries: [List of relevant event summaries $\mathbf{S}_{\theta_{k}}$ retrieved based on $q$]
Guidelines for analysis
  • General questions about overall video content (likely classified as SUMMARY) should be attempted directly from summaries.
  • Questions about specific visual details (VIDEO), sounds/speech (AUDIO), or combined elements (VIDEO+AUDIO) often require specific analysis beyond summaries.
  • Estimate confidence based on whether summaries contain sufficient information.
Output Format Must be in one of these two structures:
1. If answerable from summaries:
ANSWER: <your detailed answer based ONLY on summaries,
if multiple choice, output one letter>
@@ 2. If requiring specific analysis (summaries insufficient):
Table 8: Visual Segment Selection Prompt (Part of $\Psi_{\text{detailed}}$ - Semantic Search)
Task Description Given a question (potentially classified as VIDEO or VIDEO+AUDIO), descriptions from consolidated memory segments ($\mathbf{T}_{k}$), and an optional element to search for, identify at most 5 relevant frame indices (within segments $m_{k}$) for answering the question.
Input Question: [user query $q$] Element to search for: [search query derived from $q$]
Frame descriptions (from relevant $\mathbf{T}_{k}$): [Numbered list of frame descriptions/captions]
Instructions
  • Return ONLY numbers separated by commas (e.g., “0,3,5,8,12”) corresponding to indices within the provided list.
  • Return at most 5 indices.
  • Do not include any other text, explanations, or spaces.
  • If fewer than 5 frames are relevant, return fewer indices.
Example good responses “0,3,5,8,12” “1,4,7”
Output Format Your response (numbers only, comma-separated):
Table 9: Audio Segment Selection Prompt (Part of $\Psi_{\text{detailed}}$ - Semantic Search)
Task Description Given a question (potentially classified as AUDIO or VIDEO+AUDIO) and transcriptions with timestamps from consolidated memory segments ($\mathbf{T}_{k}$), identify the most relevant time frames where the answer might be found.
Input Question: [user query $q$]
Transcriptions (with timestamps from relevant $\mathbf{T}_{k}$): [List of timestamped transcriptions]
Instructions
  • Analyze the transcriptions and identify segments most likely to contain the answer.
  • Return a JSON array of time frames (relative to the source video). Each frame defines a start and end time.
@@
  • Return at most 5 time frames.
  • Include a small buffer around identified time frames (e.g., $\pm$2 seconds, related to $\delta$).
  • If no relevant segments found, return “[]”.
@@
Task Description You potentially have two candidate answers to the same question $q$. One derived from Fast Retrieval ($\Phi_{\text{fast}}$, using summaries $\mathbf{\Theta}$) and another from Detailed Recall ($\Psi_{\text{detailed}}$, using detailed segments $\mathcal{M}$). Compare them and provide a final, reconciled answer $a$.
Input Question: [user query $q$]
Answer from summaries (if $\Phi_{\text{fast}}$ produced one): [direct_answer] Confidence from summaries: [direct_confidence]
Answer from detailed analysis (if $\Psi_{\text{detailed}}$ was run): [detailed_answer]
Supporting Context ($\mathcal{E}_{\text{context}}$): Sample frame captions: [Sample relevant frame captions from $r_{\text{retrieved}}$] Sample transcriptions: [Sample relevant transcriptions from $r_{\text{retrieved}}$]
Instructions
  • Compare both answers for consistency, detail, and plausibility given the context.
  • If they agree, prefer the more detailed/confident answer.
  • If they disagree, evaluate which answer is better supported by the specific context (captions, transcriptions). Prioritize the answer derived from detailed analysis ($\Psi_{\text{detailed}}$) if confidence in the summary-based answer ($\Phi_{\text{fast}}$) was low or if the detailed analysis provides clearly contradictory evidence.
  • Synthesize the best possible final answer based on this comparison.
  • For multiple choice questions, select only one final answer option.
  • Estimate the confidence in the final reconciled answer.
Output Format
@@
Task Description Based on the following retrieved context ($r_{\text{retrieved}}$) and extracted evidence ($\mathcal{E}_{\text{context}}$), please synthesize a final answer ($a$) to the question ($q$).
Input Question: [user query $q$]
Overall Video Context (e.g., relevant summaries $\mathbf{S}_{\theta_{k}}$ if available): [List of relevant high-level context]
Relevant Retrieved Content (details from $\Phi_{\text{fast}}$ or $\Psi_{\text{detailed}}$): [List of retrieved content descriptions, e.g., frame captions, transcriptions, details from $m_{k}$]
Instructions Please provide a clear and specific answer based on the integration of the overall context and the specific retrieved content. If the information is insufficient or contradictory, state that or provide the best possible reasoned answer based on the evidence.
Output should be one letter if the original question was multiple choice.
Output Format
Answer: <Synthesized answer a>